what is the point of intersection of the following pair of straight lines:
6x - 8y = -1
-4x + 8y = 1

Answer by tarungehlot(9)

add both equation
2x +0=0
x=0
put x=0 in 6x - 8y=-1
we gets
-8y=-1
y=1/8
hence point of intersection
(x,y)= (0,1/8)